# Factoriality of groupoid von Neumann algebras

### v2

@@ 6.3 Full subgroupoids and partial actions
There is a Borel embedding $\iota:Y\hookrightarrow X:x\mapsto[e,x]$ and a $G$-action on $Y$ defined by $\theta_{g}([h,x])=[gh,x]$. By our construction of the equivalence relation, for any $x\in X_{g^{-1}}$, $\theta_{g}([e,x])=[g,x]=[e,\sigma_{g}(x)]$. Thus, we have $G\ltimes Y=(G\ltimes X)|_{Y}$. Finally, note that for all $(g,x)\in G\times Y$, we have $[(g,x)]=g\cdot\iota(x)$, and thus $Y$ is $G\ltimes X$-full. ∎
@@
    Let $\mu$ be a finite measure on $X$. Let $H\leq G$ and let $Y\subset X$ be an $H$-invariant Borel subset. We say that $G\curvearrowright(X,\mu)$ is induced from $H\curvearrowright Y$ if $Y$ is $\mu$-full in $G\ltimes X$ and $G\ltimes Y=H\ltimes Y$.
@@
and observe that $F(gx)=gF(x)g^{-1}$ if $x\in X_{g^{-1}}$. As in the global case, we may fix $C\subset\Gamma\setminus\{e\}$ finite for which the set
@@
In particular, if $gCg^{-1}\neq C$, then $g(Y\cap X_{g^{-1}})\cap Y=\emptyset$. Since $Y_{g}:=g(Y\cap X_{g^{-1}})\cap Y$, it follows that $Y_{g}=\emptyset$ unless $gCg^{-1}=C$. On the other hand, if $gCg^{-1}=C$, we get the identity $Y_{g}=Y\cap X_{g}$. Set $H=\{g\in G:gCg^{-1}=C\}$. We therefore have the identity $H\ltimes Y=G\ltimes Y$. Furthermore, for any $h\in C$, and $y\in Y_{h^{-1}}$, since $F(y)=C$, we have that $(h,y)\in B$. That is, $h\cdot y=y$. Thus, $h$ is an element that acts trivially on $Y$ with finite conjugacy class.
@@
In particular, for nonsingular $\mathbb{Z}$-partial actions, the groupoid $(\mathbb{Z}\ltimes X,\mu)$ is icc if and only if $\mathbb{Z}\curvearrowright(X,\mu)$ is essentially free.
@@
Conversely, assume that there is an induced subsystem $H\curvearrowright Y$ for a cyclic group $H=\langle g\rangle\leq\mathbb{F}$. Furthermore, assume that $H\curvearrowright Y$ is not essentially free. This means that there is a non-zero $n$ for which the set $E:=\{y\in Y_{g^{-n}}\cap Y_{g^{n}}\;|\;g^{n}\cdot y=y\}$ has positive measure. The set $B:=\{g^{n}\}\times E$ is a bisection in $H\ltimes Y$ with $\Omega_{B}\subseteq\{g^{n}\}\times(Y_{g^{n}}\cap Y_{g^{-n}})$, and thus $H\ltimes Y$ is icc. Since $\mathbb{F}\ltimes X$ is induced from $H\ltimes Y$, it too is icc. ∎
@@ 6.4 Deaconu–Renault Groupoids
Let $(X,\mu)$ be a finite Borel measure space and let
@@
$\operatorname{Iso}(\mathcal{G}(\sigma))=\{(x,k-l,x)\;|\;\sigma^{k}(x)=\sigma^{l}(x)\}\;.$
@@
By our previous calculations, conditions $1$ and $2$ are equivalent. That $3$ is equivalent to $2$ follows from the fact that
$s(\bigcup_{n\neq 0}B_{n})=\{x\;|\;\text{there exists }k>l\geq 0\text{ such that }\sigma^{k}(x)\neq\sigma^{l}(x)\}\;.\qed$

### v3

@@ 6.3 Full subgroupoids and partial actions
There is a Borel embedding $\iota:Y\hookrightarrow X:x\mapsto[e,x]$ and a $G$-action on $X$ defined by $\theta_{g}([h,x])=[gh,x]$. By our construction of the equivalence relation, for any $x\in Y_{g^{-1}}$, $\theta_{g}([e,x])=[g,x]=[e,\sigma_{g}(x)]$. Thus, we have $G\ltimes Y=(G\ltimes X)|_{Y}$. Finally, note that for all $(g,x)\in G\times Y$, we have $[(g,x)]=g\cdot\iota(x)$, and thus $Y$ is $G\ltimes X$-full. ∎
@@
    Let $\mu$ be a finite measure on $X$. Let $H\leq G$ and let $Y\subset X$ be an $H$-invariant Borel subset. We say that $G\curvearrowright X$ is induced from $H\curvearrowright Y$ if $Y$ is $\mu$-full in $G\ltimes X$ and $G\ltimes Y=H\ltimes Y$.
@@
and observe that $F(gx)=gF(x)g^{-1}$ if $x\in X_{g^{-1}}$. As in the global case, we may fix $C\subset G\setminus\{e\}$ finite for which the set
@@
In particular, if $gCg^{-1}\neq C$, then $g(Y\cap X_{g^{-1}})\cap Y=\emptyset$. Since $Y_{g}:=g(Y\cap X_{g^{-1}})\cap Y$, it follows that $Y_{g}=\emptyset$ unless $gCg^{-1}=C$. On the other hand, if $gCg^{-1}=C$, we get the identity $Y_{g}=Y\cap X_{g}$. Set $H=\{g\in G:gCg^{-1}=C\}$. We therefore have the identity $H\ltimes Y=G\ltimes Y$. Furthermore, for any $h\in C$ and $y\in Y_{h^{-1}}$, since $F(y)=C$, we have that $(h,y)\in B$. That is, $h\cdot y=y$. Thus, $h$ is an element that acts trivially on $Y$ with finite conjugacy class.
@@
In particular, for nonsingular partial $\mathbb{Z}$-actions, the groupoid $(\mathbb{Z}\ltimes X,\mu)$ is icc if and only if $\mathbb{Z}\curvearrowright(X,\mu)$ is essentially free.
@@
Conversely, assume that there is an induced subsystem $H\curvearrowright Y$ for a cyclic group $H=\langle g\rangle\leq\mathbb{F}$. Furthermore, assume that $H\curvearrowright Y$ is not essentially free. This means that there is a non-zero $n$ for which the set $E:=\{y\in Y_{g^{-n}}\cap Y_{g^{n}}\;|\;g^{n}\cdot y=y\}$ has positive measure. The set $B:=\{g^{n}\}\times E$ is a bisection in $H\ltimes Y$ with $\Omega_{B}\subseteq\{g^{n}\}\times(Y_{g^{n}}\cap Y_{g^{-n}})$, and thus $H\ltimes Y$ is not icc. Since $\mathbb{F}\ltimes X$ is induced from $H\ltimes Y$, it is also not icc. ∎
@@ 6.4 Deaconu–Renault Groupoids
Let $(X,\mu)$ be a Borel probability space and let
@@
$\operatorname{Iso}(\mathcal{G}(\sigma))=\{(x,k-l,x)\;|\;\sigma^{k}(x)=\sigma^{l}(x)\}\;.$ (23)
@@
Note by equation (23), $\operatorname{Iso}(\mathcal{G}(\sigma))=\bigsqcup_{n}B_{n}$. Therefore, by our previous calculations, conditions $1$ and $2$ are equivalent. That $3$ is equivalent to $2$ follows from the fact that
$s\left(\bigcup_{n\neq 0}B_{n}\right)=\{x\;|\;\text{there exists }k>l\geq 0\text{ such that }\sigma^{k}(x)\neq\sigma^{l}(x)\}\;.\qed$